adipadron:

Karen is financing $291,875 to purchase a house. She obtained a 15/5 balloon mortgage at 5.35%. What will her balloon payment be?

Extrinix:

I suggest you use this site: Mortgage Balloon Payment Calculator

kittybasil:

Based on the formula from Finance Formulas,\[\Large{FV=PV(1+r)^{n}-P[\frac{(1+r)^{n}-1}{r}]}\]where:\[FV=\text{Future Value (balloon balance)}\]\[PV=\text{Present Value (Original Balance)}\]\(P=\text{Payment, }r=\text{rate per payment, }n=\text{number of payments}\) Now let's get to work inputting the values associated with each component. This is a 15/5 balloon loan, which means it will be calculated for 15 years but everything after 5 years must be paid in a lump sum (basically all at once). This is commonly called "amortization" as you might have seen in one of the links (blue text) that I provided previous. For the use of the second formula, we are going to solve for the amount due after the \(15^{th}\) year (unless I misread that, so you really should get this double checked) Currently the values you have are: → present value \(PV=$291,875\) → rate \(r=0.0535\) → payment recurrence \(n=180\) (Again, this is based on the assumption that we are calculating for after 15 years. Are you sure that ratio or fraction is not upside down and 5/15 instead? We should also be calculating for the principal amount, which is referred to here as "Payment" (P).

kittybasil:

Edit to my previous response: Actually, that is not a principal, my bad. What I meant was "payment" as in monthly payment. So \(P\) is what you're paying each month. That being said, there doesn't seem to be a value provided for that. Are you missing any component of the question? @adipadron